# Stability analysis of a dark energy model in Rastall gravity

### v1

@@ IV.2 Pantheon Dataset
where $c$ is the speed of light. And, the apparent magnitude can be written in terms of luminosity distance as
@@
where $M$ is the absolute magnitude. Here, a dissipation between $H_{0}$ and $M$ can be observe and since distance modulus $\mu=m-M$, therefore the formula for $\chi^{2}$ is written as
@@
where $\mu_{th}$ and $\mu_{obs}$ used for theoretical and observed distance modulus respectively.
Figure 1: Figs. a, b, and c depict the likelihood contours for $H(z)$, $Pantheon$, and $H(z)+Pantheon$ dataset respectively corresponding to $1\sigma$ and $2\sigma$ confidence regions.
IV.3 Joint Datasets
To obtain the better constraint values of the model parameter, the joint dataset is used and the results are calculated by the $\chi_{HP}^{2}$ function, which is the sum of the functions $\chi_{Hub}^{2}$ and $\chi_{Pan}^{2}$ as
@@ IV.3 Joint Datasets
The obtained values of $j$ and $\alpha$ from the above-mentioned observational datasets are placed in Table-1. According to the $\Lambda$CDM model, the present value of jerk parameter $j$ is $1$ and in our model, we obtained the best-fit value of $j$ which is close to 1 for the $Pantheon$ and joint datasets and for $H(z)$ dataset it is approximately $0.435$.
@@
As the $\Lambda$CDM model is one of the standard models in cosmology and by comparing our model with the $\Lambda$CDM model we check the consistency of our model and for this purpose, error bar plots play a significant role. In our model, we have drawn the error bar plots for Hubble datasets and Type Ia Supernovae datasets. In Fig. 2(a) and Fig. 2(b), we observe the similarity between our model and the $\Lambda$CDM model with the error bar.
@@
Figure 2: The error bar plots for $H(z)$ and Type Ia Supernovae datasets.
V Physical and Dynamical Components of the model
@@
To explore the dynamical evolution of the universe, let us start with one of the essential parameters named as deceleration parameter ($q$). The deceleration parameter has been calculated in Eq. (5) and now we plot the curves for obtained best-fit values of $j$ and $\alpha$. For all three mentioned observational datasets, we perceive that $q$ is behaving alike i.e. in Fig. 3, at early times each trajectory is in the deceleration phase, and in late times as well as at present our model shows the accelerating phase of the universe. At present, the value of $q$ is $-0.25$ for $H(z)$, $Pantheon$, and joint datasets, and the value of redshift at the time of phase transition is $0.628$ for $H(z)$ dataset, $0.277$ for $Pantheon$ dataset and $0.239$ for $H(z)+Pantheon$ datasets which is clearly noticeable in the subplot of Fig. 3.
Figure 3: Plot of $q$  vs. $z$.
@@
It is believed that after the Big Bang, the radiation-dominated era and matter-dominated era both are the major sections in the evolution of the Universe. The radiation-dominated era is required to anticipate primordial nucleosynthesis whereas in the matter-dominated era, element creation takes place. Therefore, keeping these points in mind we analyze the evolution of various physical parameters of our model.
@@
For the fixed values of $\kappa=1$ and $\lambda=0.12$, the energy density and isotropic pressure have been plotted using Eqs. (14) and (15) in Fig. 4 for the obtained best-fit values of model parameters. Fig. 4(a) illustrates the evolution of energy density from high redshift to low redshift. As expected, the energy density decreases from early to late times for all datasets. In this model, the isotropic pressure is plotted in Fig. 4(b), which highlights that in the early universe, the value of pressure is highly positive and at present as well as in the future pressure is negative for all datasets. According to the standard cosmology, the negative pressure results expanding acceleration of the cosmos. The different value of EoS parameter ($\omega$) indicates the different forms of universe for example, $\omega=0$to pressureless matter, $\omega\in(0,\frac{1}{3})$to hot matter, $\omega=\frac{1}{3}$to radiation, $\omega\in(\frac{1}{3},1)$to hard universe, $\omega=1$to stiff matter, $\omega>1$to ekpyrotic matter, $\omega\in(0,-1)$to quintessence universe, $\omega=-1$to cosmological constant and $\omega<-1$to Phantom universe. In our model, one can observe the different forms of the universe with respect to the changes in the value of $\omega$ (see Fig. 4(c)). Today and in the future, our model is a quintessence dark energy model.
Figure 4: The graphs of $\rho$, $p$ and $\omega$  vs. $z$.
@@
Besides phenomenal fields, there are some major points to explore such as the classification of singularities, casual structure, energy conditions, etc. In particular, energy conditions were originally formulated from the well-known Raychaudhuri equation. As it is valuable to explain the contribution of further geometrical terms in the stress-energy tensor such as effective pressure and effective energy density we study four energy conditions. The energy conditions are the diminution of time-like or null vector fields concerning the energy-momentum tensor and Einstein tensor obtained from Einstein field equations. Generally, four energy conditions are discussed in cosmological models namely, Weak energy condition (WEC), Null energy condition (NEC), Strong energy condition (SEC), and Dominant energy condition (DEC). These ECs are defined as: WEC: $\rho\geq 0,~{}\rho+p>0$, NEC: $\rho+p\geq 0$, SEC: $\rho+3p\geq 0$, and DEC: $\rho\geq|p|$.
Figure 5: The plots for NEC, SEC, DEC.
@@
In Fig. 5, we perceive that for all observational datasets NEC and DEC hold for the whole range of $z$, whereas SEC violates in the future.
VI Dynamical system stability
@@
In this section, we employ the dynamical system technique to assess the stability of the system. To achieve this, we initially select a set of dimensionless variables as follows,
$\Omega_{M}=\frac{\kappa\rho}{3H^{2}},\quad y=\frac{\kappa p}{H^{2}},\quad\delta=\kappa\lambda,$ (21)
@@
The variable $\Omega_{M}$ represents a time-dependent quantity, signifying the fractional energy density budget of the universe, while $y$ captures the dynamics of the corresponding pressure. We introduce a dimensionless constant, denoted as $\delta$, which can take any non-zero value. Utilizing Eq. (12) and Eq. (13), we express,
@@
Expressing Eq. (22) in terms of the dynamical variables that constrain $y$, we can represent this constraint in terms of $\Omega_{M}$ as,
@@
To analyze the dynamics of the system, we construct the autonomous system by calculating the first-order time derivative of the dynamical variable, as follows,
@@
Here, $Hdt=dN=d\log a$. We have used $j=\frac{\dddot{a}}{aH^{3}}$ as a constant parameter. This constraint allows the evaluation of the system’s dynamics using a single parameter, resulting in a 1D phase space. The equation of state and deceleration parameter can then be expressed in terms of dynamical variables as,
@@
The critical points that satisfy this constraint are the only ones considered physically viable. The system produces two critical points, which are tabulated in Tab. 2. The qualitative behavior of these points is discussed in detail.
Table 2: The critical points and its nature.
@@
Figure 6: The plots for the constrained relation Eq. (27) and EoS $\omega$, in the parametric space of $(j,\delta)$ for the fixed point $P_{1}$.
@@
    Figure 7: The plots for the constrained relation Eq. (27) and EoS parameter $\omega$, in the parametric space of $(j,\delta)$ for the fixed point $P_{2}$.
@@
From the analysis of critical points and based on their characteristic properties, we have constrained $j$ and $\delta$. One of the critical points is found to be stable, producing an accelerating solution that represents the present epoch of the universe. To further analyze the system’s dynamics, we evaluate the evolution of cosmological parameters, i.e., $(\Omega_{M},\omega,h\equiv H/100)$ against $N=\log a$ in Fig. 8, spanning from the past epoch to the future epoch. We select benchmark points $(j=0.8,\delta=0.04)$ to ensure the existence of both fixed points simultaneously. In the past epoch, the fractional energy density dominates, and the corresponding equation of state is near 00, indicating a matter-dominated phase of the universe. As the system evolves toward the present epoch, the equation of state increases towards negative values. In the future epoch, $\omega$ saturates to $-0.70$, and energy density saturates to $0.90$. Hence, the system exhibits quintessence-like behavior in the late-time epoch and produces matter characteristics in the past epoch of the universe. We have also depicted the evolution of the Hubble parameter $h(N)=H(N)/100$ using
@@
VII Conclusion
In this work, we study Rastall’s theory of gravity, which is based on a variational principle. In this theory, we consider a non-minimal coupling between geometry and matter fields and discuss the recent theoretical developments in alternative theories of general relativity. This work is framed for the flat, isotropic, and homogeneous universe in the FLRW metric, where we have obtained the EFEs with the calculated value of the Hubble parameter in terms of the jerk parameter. Furthermore, we examine the essential physical and dynamical parameters one by one. We evaluate the best-fitted values of model parameters $\alpha$ and $j$ by the MCMC method, where we use the Hubble dataset, $Pantheon$ dataset, and their joint dataset.
The trajectories of the deceleration parameter for these best-fit values indicate the phase transition from deceleration to acceleration. At present, our model shows the accelerating universe.
Also, if we look at the behavior of energy density, it is monotonically decreased from high redshift to low redshift for all observational datasets. The isotropic pressure $p$ is positive in the early universe and holds a negative value at present as well as in the future, which indicates the presence of repulsive force and an accelerating expansion of the universe. According to the results of the EoS parameter, our model is a quintessence dark energy model at present. In Fig. 4, $\omega$ depicts the different dominating eras of the universe.
In this model, NEC and DEC both hold good but SEC violates which shows the existence of exotic matter in the universe. Hence, finally, we conclude that our model which is studied in Rastall gravity is a quintessence dark energy model.
Finally, we conducted a dynamical stability analysis of the system, treating the jerk parameter $j$ as a constant quantity, which imposes a vital constraint on the system. Our findings revealed a 1D phase space, with the system yielding two critical points. One critical point results in a stable accelerating solution in the future epoch, resembling the behavior of a quintessential dark energy model. The other critical point exhibits non-accelerating characteristics, similar to dark matter fluid, and stabilizes at the past epoch. Thus, the qualitative behavior of the system demonstrates that Rastall’s theory of gravity can consistently account for different phases, namely the matter-dominated and accelerating epochs of the universe.
@@ Acknowledgements
J. K. Singh wishes to thank Joao Luis Rosa, Institute of Physics, University of Tartu, W. Ostwaldi 1, 50411 Tartu, Estonia for fruitful discussions.
Data Availability Statement No new data were created or analyzed in this study.
@@ Appendix A Dynamical system in 1D
Consider the autonomous equation in 1D is:

### v2

@@ V.2 Pantheon Dataset
where $c$ is the speed of light. And the apparent magnitude can be written in terms of luminosity distance as
@@
where $M$ is the absolute magnitude. Here, a dissipation between $H_{0}$ and $M$ can be observed, and since the distance modulus $\mu=m-M$, the formula for $\chi^{2}$ is written as
@@
where $\mu_{th}$ and $\mu_{obs}$ are used for theoretical and observed distance moduli, respectively.
Figure 1: Depiction of likelihood contours for the $H(z)$, $Pantheon$, and $H(z)+Pantheon$ datasets within the $1\sigma$ and $2\sigma$ confidence levels.
V.3 Joint Dataset
To obtain better constraint values of the model parameters, we apply a combined dataset. The results are computed using the $\chi_{HP}^{2}$ function, which is the sum of the $\chi_{Hub}^{2}$ and $\chi_{Pan}^{2}$ functions, as shown below
@@ V.3 Joint Dataset
The tabulated results in Table-1 showcase the acquired values of $j$ and $\alpha$ based on the aforementioned observational datasets. While the $\Lambda$CDM model has the present value of the jerk parameter $j$ as $1$, our model obtains the best-fit value of $j$ close to $1$ for the $Pantheon$ and joint datasets, and for the $H(z)$ dataset, it is approximately $0.437$.
@@
As the $\Lambda$CDM model is one of the standard models in cosmology, by comparing our model with the $\Lambda$CDM model, we check the consistency of our model, and for this purpose, error bar plots play a significant role. In our model, we have drawn the error bar plots for Hubble datasets and Type Ia Supernovae datasets. In Fig. 2(a) and Fig. 2(b), we observe the similarity between our model and the $\Lambda$CDM model.
@@
Figure 2: Visual display of error bars for $H(z)$ and Type Ia Supernovae datasets.
VI Physical and Dynamical Components of the model
@@
To explore the dynamical progression of the universe, let us start with one of the essential parameters named as the deceleration parameter ($q$). The deceleration parameter has been calculated in Eq. (10), and now we plot the curves for the obtained best-fit values of $j$ and $\alpha$. For all three mentioned observational datasets, we perceive that $q$ is behaving alike, i.e., in Fig. 3, at early times each trajectory is in the deceleration phase, and in late times as well as at present, our model shows the accelerating phase of the universe. At present, the value of $q$ is $-0.25$ for $H(z)$, $Pantheon$, and joint datasets, and the value of redshift at the time of phase transition is $0.628$ for $H(z)$ dataset, $0.277$ for $Pantheon$ dataset, and $0.239$ for $H(z)+Pantheon$ dataset, which is clearly noticeable in the subplot of Fig. 3.
Figure 3: Plot of $q$  vs. $z$ for the mentioned observational datasets.
@@
It is widely accepted that following the Big Bang, the evolution of the universe is marked by two significant epochs: the radiation-dominated era and the matter-dominated era. The radiation-dominated era is required to anticipate primordial nucleosynthesis, whereas in the matter-dominated era, element creation takes place. Therefore, keeping these points in mind, we analyze the evolution of diverse physical parameters within our model.
@@
For the fixed values $\kappa=1$ and $\lambda=0.12$, the energy density and isotropic pressure have been plotted using Eqs. (19) and (20) in Fig. 4 for the obtained best-fit values of model parameters. In Fig. 4(a), we depict the change in energy density across varying redshifts. As expected, the energy density decreases from early to late times for all datasets. In this model, the isotropic pressure is plotted in Fig. 4(b), which highlights that in the early universe, the value of pressure is highly positive, and at present as well as in the future, pressure is negative for all datasets. According to standard cosmology, the negative pressure reveals the expanding acceleration of the cosmos. The different value of the EoS parameter ($\omega$) indicates the different forms of the universe, for example, $\omega=0$to pressureless matter, $\omega\in(0,\frac{1}{3})$to hot matter, $\omega=\frac{1}{3}$to radiation, $\omega\in(\frac{1}{3},1)$to hard universe, $\omega=1$to stiff matter, $\omega>1$to ekpyrotic matter, $\omega\in(-1,-\frac{1}{3})$to quintessence universe, $\omega=-1$to cosmological constant, and $\omega<-1$to phantom universe. In our model, one can observe the different forms of the universe with respect to the changes in the value of $\omega$ (see Fig. 4(c)). Today and in the future, our model aligns with a quintessence dark energy model.
Figure 4: Visual display of $\rho$, $p$, and $\omega$  vs. $z$ for the mentioned observational datasets.
@@
In addition to the fascinating fields of study, there exist several major points for exploration, including the characterization of singularities, causal structure, and energy conditions. Originating from the renowned Raychaudhuri equation, energy conditions play a pivotal role in explaining the contribution of further geometrical terms in the stress-energy tensor, such as effective pressure and effective energy density. We study four energy conditions, which are the diminution of time-like or null vector fields concerning the energy-momentum tensor and the Einstein tensor obtained from Einstein field equations. Generally, four energy conditions are discussed in cosmological models namely, Weak energy condition (WEC), Null energy condition (NEC), Strong energy condition (SEC), and Dominant energy condition (DEC). These ECs are defined as: WEC: $\rho\geq 0,~{}\rho+p>0$, NEC: $\rho+p\geq 0$, SEC: $\rho+3p\geq 0$, and DEC: $\rho\geq|p|$.
Figure 5: The plots for NEC, SEC, and DEC for the mentioned observational datasets.
@@
In Fig. 5, our probe reveals that for all observational datasets, NEC and DEC hold for the whole range of $z$, whereas SEC violates in the future.
VII Dynamical System Stability
@@
In this section, we employ the dynamical system technique to assess the stability of the system. To achieve this, we initially select a set of dimensionless variables as follows
$\Omega_{M}=\frac{\kappa\rho}{3H^{2}},\quad y=\frac{\kappa p}{H^{2}},\quad\delta=\kappa\lambda.$ (26)
@@
The variable $\Omega_{M}$ represents a time-dependent quantity, signifying the fractional energy density budget of the universe, while $y$ captures the dynamics of the corresponding pressure. We introduce a dimensionless constant, denoted as $\delta$, which can take any non-zero value. Utilizing Eq. (17) and Eq. (18), we express
@@
Expressing Eq. (27) in terms of the dynamical variables that constrain $y$, we can represent this constraint in terms of $\Omega_{M}$ as
@@
To analyze the dynamics of the system, we construct the autonomous system by calculating the first-order time derivative of the dynamical variable, as follows
@@
Here, $Hdt=dN=d\log a$. We have used $j=\frac{\dddot{a}}{aH^{3}}$ as a constant parameter. This constraint allows the evaluation of the system’s dynamics using a single parameter, resulting in a 1D phase space. The equation of state and deceleration parameter can then be expressed in terms of dynamical variables as
@@
The critical points that satisfy this constraint are the only ones that are considered physically viable. The system produces two critical points, which are tabulated in Table-2. The qualitative behavior of these points is discussed in detail.
Table 2: The critical points and their nature
@@
Figure 6: The plots for the constrained relation Eq. (32) and EoS $\omega$ in the parametric space of $(j,\delta)$ for the fixed point $P_{1}$.
@@
    Figure 7: The plots for the constrained relation Eq. (32) and EoS parameter $\omega$ in the parametric space of $(j,\delta)$ for the fixed point $P_{2}$.
@@
From the analysis of critical points and based on their characteristic properties, we have constrained $j$ and $\delta$. One of the critical points is found to be stable, producing an accelerating solution that represents the present epoch of the universe. To further analyze the system’s dynamics, we evaluate the evolution of cosmological parameters, i.e., $(\Omega_{M},\omega,h\equiv H/100)$ against $N=\log a$ in Fig. 8, spanning from the past epoch to the future epoch. We select benchmark points $(j=0.8,\delta=0.04)$ to ensure the existence of both fixed points simultaneously. In the past epoch, the fractional energy density dominates, and the corresponding equation of state is near zero, indicating a matter-dominated phase of the universe. As the system evolves toward the present epoch, the equation of state increases toward negative values. In the future epoch, $\omega$ saturates to $-0.70$, and energy density saturates to $0.90$. Hence, the system exhibits quintessence-like behavior in the late-time epoch and produces matter characteristics in the past epoch of the universe. We have also depicted the evolution of the Hubble parameter $h(N)=H(N)/100$ using
@@
VIII Conclusions
In this work, we study Rastall’s theory of gravity, which is based on a variational principle. In this theory, we consider a non-minimal coupling between geometry and matter fields and discuss the recent theoretical developments in alternative theories of general relativity. This work is framed for the flat, isotropic, and homogeneous universe in the FLRW metric, where we have obtained the EFEs with the calculated value of the Hubble parameter in terms of the jerk parameter. Furthermore, we examine the essential physical and dynamical parameters one by one. We evaluate the best-fitted values of model parameters $\alpha$ and $j$ by the MCMC method, where we use the Hubble dataset, the $Pantheon$ dataset, and their joint dataset.
The trajectories of the deceleration parameter for these best-fit values indicate the phase transition from deceleration to acceleration. At present, our model shows an accelerating universe.
Also, if we look at the behavior of energy density, it has monotonically decreased from high redshift to low redshift for all observational datasets. The isotropic pressure $p$ is positive in the early universe and holds a negative value at present as well as in the future, which indicates the presence of repulsive force and an accelerating expansion of the universe. According to the results of the EoS parameter, our model is a quintessence dark energy model at present. In Fig. 4, $\omega$ depicts the different dominating eras of the universe.
In this model, NEC and DEC both hold good, but SEC violates, which shows the existence of exotic matter in the universe. Hence, finally, we conclude that our model, which is studied in Rastall gravity, is a quintessence dark energy model.
Finally, we conducted a dynamical stability analysis of the system, treating the jerk parameter $j$ as a constant quantity, which imposes a vital constraint on the system. Our findings revealed a 1D phase space, with the system yielding two critical points. One critical point results in a stable, accelerating solution in the future epoch, resembling the behavior of a quintessential dark energy model. The other critical point exhibits non-accelerating characteristics, similar to dark matter fluid, and stabilizes at the past epoch. Thus, the qualitative behavior of the system demonstrates that Rastall’s theory of gravity can consistently account for different phases, namely the matter-dominated and accelerating epochs of the universe.
@@ Acknowledgements
J. K. Singh wishes to thank Prof. Joao Luis Rosa, Institute of Physics, University of Tartu, W. Ostwaldi 1, 50411 Tartu, Estonia, for fruitful discussions.
Data Availability Statement
No new data were created or analyzed in this study.
Code Availability Statement
No new code was created in this study.
@@ Appendix A Dynamical system in 1D
Consider the autonomous equation in 1D as,